# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ Abstract
In this paper, we prove directional convergence of network parameters of fixed width leaky ReLU two-layer neural networks optimized by gradient descent with exponential loss, which was previously only known for gradient flow. By a careful analysis of the convergent direction, we establish sufficient conditions of benign overfitting and discover a new phase transition in the test error bound. All of these results hold beyond the nearly orthogonal data setting which was studied in prior works. As an application, we demonstrate that benign overfitting occurs with high probability in sub-Gaussian mixture models.
@@ 1 Introduction
In this paper, we investigate benign overfitting in two-layer leaky ReLU neural networks with fixed width for binary classification whose network parameters are optimized by gradient descent with exponential loss.
Our main results are summarized as follows:
  1. 1.
    Directional Convergence: we show that the network parameter converges in direction if the parameter is optimized by gradient descent with exponential loss. In the existing literature, the directional convergence was proven only for gradient flow by Ji and Telgarsky (2020). For gradient descent, only weaker forms of implicit bias were shown (Lyu and Li (2020); Frei et al. (2023b); Kou et al. (2023b)). Moreover, the analysis in Frei et al. (2023b); Kou et al. (2023b) is limited to the rather specialized setting of nearly orthogonal data, i.e. the predictors $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,n}\subset\mathbb{R}^{p}$ satisfy $\|\boldsymbol{x}_{i}\|^{2}\gg\max_{i\neq k}|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle|$.
  2. 2.
    Benign Overfitting: as an application of the directional convergence, we establish sufficient conditions for benign overfitting by analyzing the property of convergent direction. Due to the difficulty in establishing directional convergence, prior works for leaky ReLU networks were limited to either gradient flow (Frei et al. (2023a)), smoothed approximation (Frei et al. (2022)), or nearly orthogonal data regime (Xu and Gu (2023)). Taking advantage of a careful analysis of the convergent direction, our work establishes sufficient conditions for benign overfitting beyond previously analyzed regimes, which reveals a new phase transition in the test error bound.
  3. 3.
    Beyond NTK: both the directional convergence and benign overfitting are shown for fixed width two-layer neural networks, thus our result is beyond the so called neural tangent kernel (NTK) regime or the lazy training regime (Jacot et al. (2018); Arora et al. (2019)).
@@ 2 Related Work
Most prior works on the implicit bias of (leaky) ReLU neural networks for binary classification have considered gradient flow. For gradient flow, Lyu and Li (2020) showed that any limit point of the network parameter of homogeneous neural networks trained with exponential type loss is a KKT point of a constrained optimization problem and that margin maximization occurs. Ji and Telgarsky (2020) further established directional convergence of the parameter of homogeneous neural networks trained with the exponential type loss. Taking advantage of the directional convergence of gradient flow, Lyu et al. (2021); Bui Thi Mai and Lampert (2021); Sarussi et al. (2021); Frei et al. (2023b); Min et al. (2024) obtained detailed characterizations of the convergent direction of the parameters of (leaky) ReLU two-layer neural networks under further assumptions on on the training data.
In contrast to gradient flow, directional convergence of network parameters was not proven for gradient descent with exponential type loss, even under strong assumptions. As a result, only weaker forms of implicit bias have been obtained for gradient descent. Lyu and Li (2020) showed that margin maximization also occurs for the network parameters of a homogeneous network trained by gradient descent, but their assumptions include smoothness of the network, and hence rule out (leaky) ReLU networks. Frei et al. (2023b) showed that the stable rank of network parameters stays bounded by a constant for two-layer neural networks, but their result is limited to the smoothed approximation of leaky ReLU. Kou et al. (2023b) proved similar results for (leaky) ReLU two-layer neural networks assuming that the number of positive and negative neurons are equal. Moreover, Frei et al. (2023b); Kou et al. (2023b) are limited to the rather specialized scenario of nearly orthogonal data.
Benign overfitting in neural networks for binary classification.
As noted earlier, benign overfitting in neural networks remains poorly understood. In fact, very few theoretical results are established beyond the NTK regime. Cao et al. (2022) studied benign overfitting for two-layer convolutional neural networks with polynomial ReLU activation ($\mathrm{ReLU}^{q},q>2$). Their results were strengthened by Kou et al. (2023a) by relaxing the polynomial ReLU assumption. Both Cao et al. (2022) and Kou et al. (2023a) are not applicable to our setting since our work considers fully connected neural networks.
More closely related work is Frei et al. (2022), which studied benign overfitting in smoothed leaky ReLU two-layer neural networks on mixture data. Their work was extended by Xu and Gu (2023) to non-smooth leaky ReLU like activation. While Xu and Gu (2023)’s work is not limited to (leaky) ReLU, their work is limited to the nearly orthogonal data regime and only holds for isotropic mixture with bounded log-Sobolev constant, which is a stronger assumption than sub-Gaussian mixture. Our work establishes benign overfitting beyond the nearly orthogonal data regime and can be applied to a wider class of mixtures than prior works.
A concise summary of related prior works is given in Table 1.
@@
Gradient Descent NA* NA Frei et al. (2022)**, NA
@@
* Some weaker forms of implicit bias were studied by Frei et al. (2023b); Kou et al. (2023b).
@@ 3 Preliminaries
where $\alpha>0$ is a fixed constant and we assume $\phi^{\prime}(0)$ is any number in $[\gamma,1]$. The second layer weights $a_{j}$ are fixed throughout the training. Fixing the second layer weights $a_{j}\in\{\pm 1/\sqrt{m}\}$ and only optimizing the first layer weights $\boldsymbol{w}_{j}$ is standard in the existing literature, see e.g. Arora et al. (2019); Cao et al. (2022); Frei et al. (2022); Xu and Gu (2023); Kou et al. (2023b). We do not consider random initialization of $W^{(0)}$ since only the magnitude of $W^{(0)}$ plays a role in our analysis.
@@
    The observations consist of $n$ i.i.d copies $(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n$ of the pair $(\boldsymbol{x},y)$. Here for a random variable $y\in\{-1,1\}$ satisfying $P(y=1)=P(y=-1)=1/2$, a random vector $\boldsymbol{z}\in\mathbb{R}^{p}$ independent of $y$, and a deterministic $\boldsymbol{\mu}\in\mathbb{R}^{p}$, we have
@@
As a special case of the model (M), we consider sub-Gaussian mixture models:
@@
    Suppose in model (M) $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$, where $\boldsymbol{\xi}\in\mathbb{R}^{p}$ has independent entries $\xi_{j}$ that have mean zero, unit variance, and satisfy $\|\xi_{j}\|_{\psi_{2}}\leq L$ for all $j=1,\ldots,p$.
The mixture models (M) and (sG) have been studied for linear classifiers by Chatterji and Long (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025) and for two-layer neural networks with $\Sigma=I_{p}$ by Frei et al. (2022); Xu and Gu (2023).
@@
4 Directional Convergence
We start from our main results on directional convergence of leaky ReLU two-layer neural networks.
The key to show the directional convergence of the gradient descent iterates (2) is activation of neurons. We say that $j$-th neuron $\phi(\langle\cdot,\boldsymbol{w}_{j}\rangle)$ is activated if $a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle)>0$ holds for any $i\in I$, which can be equivalently written as $a_{j}y_{i}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle>0$ for any $i\in I$. We will later establish conditions under which $a_{j}y_{i}\langle\boldsymbol{x}_{i},w_{j}^{(t)}\rangle$ monotonically increases as the gradient descent progresses.
Our first result shows that if all the neurons stay activated after some step of gradient descent (2), directional convergence occurs towards the maximum margin vector $\hat{W}$ in a transformed sample space, i.e. $\hat{W}$ is the solution to the following optimization problem:
where $\boldsymbol{\tilde{x}}_{i}^{\top}=(\boldsymbol{\tilde{x}}_{i,1}^{\top},\boldsymbol{\tilde{x}}_{i,2}^{\top},\ldots,\boldsymbol{\tilde{x}}_{i,m}^{\top})\in\mathbb{R}^{1\times mp}$ with $\boldsymbol{\tilde{x}}_{i,j}=a_{j}\zeta(a_{j}y_{i})\boldsymbol{x}_{i}$. Letting $\sigma_{max}(\tilde{X})$ be the maximum singular value of $\tilde{X}\in\mathbb{R}^{n\times mp}$ whose rows consist of $\boldsymbol{\tilde{x}}_{i}^{\top}$’s, we have
Theorem 4.1.
Suppose all the neurons are activated for $t\geq T$ and $\alpha\sigma_{max}^{2}(\tilde{X})$ <2. Then, the gradient descent iterate (2) satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$ and $W^{(t)}$ converges in direction to $\hat{W}$.
Lastly, the network $f(\cdot;\hat{W})$ has the linear decision boundary defined by $\boldsymbol{\bar{w}}:=\frac{|J_{+}|}{\sqrt{m}}\boldsymbol{w}_{+}-\frac{|J_{-}|}{\sqrt{m}}\boldsymbol{w}_{-}$, i.e. $\mathrm{sign}(f(\boldsymbol{x}:\hat{W}))=\mathrm{sign}(\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle)$ for any $\boldsymbol{x}\in\mathbb{R}^{p}$.
Previously, the characterizations of the convergent direction in Theorem 4.1 were only obtained for gradient flow trained on nearly orthogonal data by Frei et al. (2023b). The proof of Theorem 4.1 is given in Appendix A.1.
To demonstrate that Theorem 4.1 can guarantee directional convergence of the network parameter $W^{(t)}$ beyond the nearly orthogonal regime, we verify the conditions of the theorem in model (M). Specifically, we find conditions which ensure that the neurons stay activated and $\alpha\sigma_{max}^{2}(\tilde{X})<2$.
In the following, we let $R_{max}(\boldsymbol{z}):=\max_{i}\|\boldsymbol{z}_{i}\|$ and $R_{min}(\boldsymbol{z}):=\min_{i}\|\boldsymbol{z}_{i}\|$, and define events that capture deterministic conditions the training data need to satisfy for the directional convergence:
@@ 4 Directional Convergence
In addition, we denote by $\sigma:=\max_{j}\|\boldsymbol{w}_{j}\|$ the size of initialization. It is also convenient to let $\rho:=\sigma\sqrt{m(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}}$ to make the assumptions of our main results concise.
We present the sufficient conditions of directional convergence for the following two cases:
Case 1: $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$.
In this case, we make the following assumptions:
@@
The intuition behind these assumption is as follows: It can be shown that $\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle\geq 0$ for any $i\neq k$ under $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$. Assumption 4.2 ensures the step size sufficiently larger than initialization so that neuron activation occurs immediately after one step of (2). Then, Assumption 4.3 ensures $\alpha\tilde{\sigma}_{max}^{2}(\tilde{X})<2$ so that the directional convergence holds.
Theorem 4.4 (Directional Convergence, Case 1).
Suppose event $E$ holds with Assumptions 4.2 and 4.3 with $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ and $\theta_{2}<1$. Then, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
We note that there is no upper bound to the signal $\|\boldsymbol{\mu}\|$. Therefore, Theorem 4.4 allows non-nearly-orthogonal data. The proof of Theorem 4.4 is given in Appendix A.2.
Case 2: Weak Signal Regime.
To deal with the case when the signal is not strong enough so that $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle,k\neq i$ could be negative, we consider the following assumptions:
@@
$16\alpha R_{max}^{2}(\boldsymbol{z})\exp(\rho)<$ 1 .
Assumption 4.5 and 4.7 allow us to carefully control $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle$. Small $\varepsilon_{1},\varepsilon_{3}$ ensure that the training data are nearly orthogonal, and small $\varepsilon_{2}$ requires sufficiently small initialization relatively to step size. By taking $\boldsymbol{\mu}=0$, this case includes the nearly orthogonal data regime studied in prior works (Frei et al. (2023b); Kou et al. (2023b)). The other assumptions are analogous to Case 1.
With these assumptions, we have
Theorem 4.9 (Directional Convergence for Weak Signal Regime, Case 2).
Suppose event $E$ holds under Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2},i=1,3$, and $\theta_{2}\leq\tfrac{1}{2}$, where $C_{w}$ is a constant defined by
$C_{w}:=\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}\exp(1).$
Then, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
The proof of Theorem 4.9 is given in Appendix A.3.
We note that Case 1 was not studied in prior works while Case 2 essentially corresponds to the near orthogonal regime studied in Frei et al. (2023b); Kou et al. (2023b). Therefore, theorems in this section not only strengthen existing results but also establish directional convergence in a regime not studied in the existing literature.

### v2

@@ Abstract
In this paper, we study benign overfitting of fixed width leaky ReLU two-layer neural network classifiers trained on mixture data via gradient descent. We provide both, upper and lower classification error bounds, and discover a phase transition in the bound as a function of signal strength. The lower bound leads to a characterization of cases when benign overfitting provably fails even if directional convergence occurs. Our analysis allows us to considerably relax the distributional assumptions that are made in existing work on benign overfitting of leaky ReLU two-layer neural network classifiers. We can allow for non-sub-Gaussian data and do not require near orthogonality. Our results are derived by establishing directional convergence of the network parameters and studying classification error bounds for the convergent direction. Previously, directional convergence in (leaky) ReLU neural networks was established only for gradient flow. By first establishing directional convergence, we are able to study benign overfitting of fixed width leaky ReLU two-layer neural network classifiers in a much wider range of scenarios than was done before.
@@ 1 Introduction
In this work, we investigate benign overfitting in two-layer leaky ReLU neural networks with fixed width for binary classification whose network parameters are trained on mixture data by gradient descent with exponential loss. We show that benign overfitting occurs in a much wider scenarios than previously known.
Our results are derived by establishing directional convergence of the network parameters with precise characterization of the convergent direction and by obtaining a classification error bound of that direction. Due to the difficulty in establishing directional convergence of gradient descent for ReLU type networks, prior work for leaky ReLU networks were limited to either gradient flow (Frei et al. (2023a)), smoothed approximation of leaky ReLU (Frei et al. (2022)), or nearly orthogonal data regime (Xu and Gu (2023)), i.e. the predictors $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,n}\subset\mathbb{R}^{p}$ satisfy $\|\boldsymbol{x}_{i}\|^{2}\gg\max_{i\neq k}|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle|$.
The precise characterization of convergent direction not only allows us to prove benign overfitting beyond nearly orthogonal data regime but also leads to a novel lower bound of the classification error in the case of Gaussian mixture. The lower bound shows our bound is tight and allows us to characterize cases when benign overfitting provably fails even if directional convergence occurs.
Moreover, sufficient conditions of directional convergence and classification error bound are obtained in a deterministic manner, which allows us to show that benign overfitting occurs with high probability on polynomially tailed mixture models while the prior results are limited to sub-Gaussian mixture.
@@ 2 Related Work
Benign overfitting in neural networks for binary classification.
As noted earlier, benign overfitting in neural networks remains poorly understood. In fact, very few theoretical results are established beyond the so called neural tangent kernel (NTK) regime. Cao et al. (2022) studied benign overfitting for two-layer convolutional neural networks with polynomial ReLU activation ($\mathrm{ReLU}^{q},q>2$). Their results were strengthened by Kou et al. (2023a) by relaxing the polynomial ReLU assumption. Both Cao et al. (2022) and Kou et al. (2023a) are not applicable to our setting since our work considers fully connected neural networks.
More closely related work is Frei et al. (2022), which studied benign overfitting in smoothed leaky ReLU two-layer neural networks on mixture data. Their work was extended by Xu and Gu (2023) to non-smooth leaky ReLU like activation. While Xu and Gu (2023)’s work is not limited to (leaky) ReLU, their work is limited to the nearly orthogonal data regime and only holds for isotropic mixture with bounded log-Sobolev constant, which is a stronger assumption than sub-Gaussian mixture. Our work establishes benign overfitting beyond the nearly orthogonal data regime and can be applied to a wider class of mixtures than the prior work such as polynomially tailed mixture. Moreover, we also establish a novel lower bound of the classification error in the case of Gaussian mixture, which shows tightness of our result and helps characterize cases when benign overfitting provably fails. Even with gradient flow, benign overfitting in leaky ReLU neural networks is limited to nearly orthogonal data setting and sub-Gaussian mixture model (Frei et al. (2023a)).
Lastly, our result is shown for fixed width two-layer neural networks, which relaxes Xu and Gu (2023) who required $m$ to grow as $n$, and thus is also beyond the NTK regime or the lazy training regime (Jacot et al. (2018); Arora et al. (2019)).
@@
Most prior work on the implicit bias of (leaky) ReLU neural networks for binary classification has considered gradient flow. For gradient flow, Lyu and Li (2020) showed that any limit point of the network parameter of homogeneous neural networks trained with exponential type loss is a KKT point of a constrained optimization problem and that margin maximization occurs. Ji and Telgarsky (2020) further established directional convergence of the parameter of homogeneous neural networks trained with the exponential type loss. Taking advantage of the directional convergence of gradient flow, Lyu et al. (2021); Bui Thi Mai and Lampert (2021); Sarussi et al. (2021); Frei et al. (2023b); Min et al. (2024) obtained detailed characterizations of the convergent direction of the parameters of (leaky) ReLU two-layer neural networks under further assumptions on the training data.
In contrast to gradient flow, directional convergence of network parameters for ReLU type neural networks was not proven for gradient descent with exponential type loss, even under strong assumptions. As a result, only weaker forms of implicit bias have been obtained for gradient descent. Lyu and Li (2020) showed that margin maximization also occurs for the network parameters of a homogeneous network trained by gradient descent, but their assumptions include smoothness of the network, and hence rule out (leaky) ReLU networks. Frei et al. (2023b) showed that the stable rank of network parameters stays bounded by a constant for two-layer neural networks, but their result is limited to the smoothed approximation of leaky ReLU. Kou et al. (2023b) proved similar results for (leaky) ReLU two-layer neural networks assuming that the number of positive and negative neurons are equal. Moreover, Frei et al. (2023b); Kou et al. (2023b) are limited to the rather specialized scenario of nearly orthogonal data regime. Cai et al. (2025) showed directional convergence of gradient descent, but their result is not applicable to ReLU type neural networks since their result requires networks to be twice differentiable with respect to the parameters. Schechtman and Schreuder (2025) extended Lyu and Li (2020) by removing twice differentiability but their result does not guarantee directional convergence and assumes that training is already in the late-stage.
To the best of our knowledge, our work is the first to show directional convergence of gradient descent in ReLU type neural networks with precise characterization of the convergent direction.
A concise summary of related prior work is given in Table 1.
@@
Gradient Descent NA* Frei et al. (2022)**, NA
@@
* Cai et al. (2025) showed directional convergence assuming twice differentiability
@@ 3 Preliminaries
where $\alpha>0$ is a fixed constant and we assume $\phi^{\prime}(0)$ is any number in $\gamma,1]$. The second layer weights $a_{j}$ are fixed throughout the training. Fixing the second layer weights $a_{j}\in\{\pm 1/\sqrt{m}\}$ and only optimizing the first layer weights $\boldsymbol{w}_{j}$ is standard in the existing literature, see e.g. Arora et al. (2019); Cao et al. (2022); Frei et al. (2022); Xu and Gu (2023); Kou et al. (2023b). We do not consider random initialization of $W^{(0)}$ since only the magnitude of $W^{(0)}$ plays a role in our analysis.
@@
    The observations consist of $n$ i.i.d copies $(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n$ of the pair $(\boldsymbol{x},y)$. Here for a random variable $y\in\{-1,1\}$ satisfying P(y=1)=P(y=1)=1/2P(y=1)=P(y=-1)=1/2, a random vector $\boldsymbol{z}\in\mathbb{R}^{p}$ independent of $y$, and a deterministic $\boldsymbol{\mu}\in\mathbb{R}^{p}$, we have
@@
As special case of the model (M), we consider sub-Gaussian mixture model (sG) and polynomially tailed mixture model (PM):
@@
    Suppose in model (M) $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$, where $\boldsymbol{\xi}\in\mathbb{R}^{p}$ has independent entries $\xi_{j}$ that have mean zero, unit variance, and satisfy $\|\xi_{j}\|_{\psi_{2}}\leq L$ for all $j=1,\ldots,p$ for $L<\infty$,
  1. (PM)
    Suppose in model (M) $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$, where $\boldsymbol{\xi}\in\mathbb{R}^{p}$ has independent entries $\xi_{j}$ that have mean zero, unit variance, and satisfy $\mathbb{E}|\xi_{j}|^{r}\leq K$ for all $j=1,\ldots,p$ for $r\in(2,4]$ and $K<\infty$.
The mixture model (sG) have been studied for linear classifiers by Chatterji and Long (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025) and for two-layer neural networks with $\Sigma=I_{p}$ by Frei et al. (2022); Xu and Gu (2023). The model (PM) has been studied for linear classifiers by Hashimoto et al. (2025).
@@
4 Directional Convergence of Gradient Descent
We start from our main result on directional convergence of gradient descent in leaky ReLU two-layer neural networks trained on mixture data. Our result provides precise characterization of the convergent direction, which is a key for our analysis on benign overfitting.
In the following, we let $R_{max}(\boldsymbol{z}):=\max_{i}\|\boldsymbol{z}_{i}\|$ and $R_{min}(\boldsymbol{z}):=\min_{i}\|\boldsymbol{z}_{i}\|$, and define an event that capture deterministic conditions the training data need to satisfy for the directional convergence:
@@ 4 Directional Convergence of Gradient Descent
In addition, we denote by $\sigma:=\max_{j}\|\boldsymbol{w}_{j}\|$ the size of initialization. It is also convenient to let $\rho:=\sigma\sqrt{m(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}}$ to keep the assumptions of our main result concise.
We present sufficient conditions of directional convergence for the following two cases:
Case 1:
This is the case when $\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle\geq 0$ is guaranteed for any $i\neq k$. In this case, we make the following assumptions:
Assumption 4.1 (Positive Correlation).
$\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z}).$
@@
The intuition behind these assumptions is as follows: It can be shown that Assumption 4.1 ensures $\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle\geq 0$ for any $i\neq k$. Assumption 4.2 ensures the step size sufficiently larger than initialization so that neuron activation occurs immediately after one step of (2). Assumption 4.3 ensures the step size is small enough for directional convergence to hold.
Case 2:
This is when $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle,k\neq i$ could be negative. The assumptions we make are:
@@
$6\alpha R_{max}^{2}(\boldsymbol{z})\exp(\rho)<1.$
Assumption 4.4 and 4.6 allow us to carefully control $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle$. Small $\varepsilon_{1},\varepsilon_{3}$ ensure that the training data are nearly orthogonal. By taking $\boldsymbol{\mu}=0$, this case includes the nearly orthogonal data regime studied in the prior work (Frei et al. (2023b); Kou et al. (2023b)). The other assumptions are analogous to Case 1.
We are now ready to state our main result on directional convergence, which not only establishes directional convergence but also provides detailed characterization of the convergent direction:
Theorem 4.8 (Directional Convergence on Mixture Data).
Suppose event $E$ holds under one of the following conditions:
  1. (i)
    Assumptions 4.1, 4.2 and 4.3 with $\theta_{2}<1$,
  2. (ii)
    Assumption 4.4, 4.5, 4.6, and 4.7 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2},i=1,3$, and $\theta_{2}\leq\tfrac{1}{2}$, where $C_{w}:=24\exp(1)\gamma^{-2}R_{max}^{2}(\boldsymbol{z})R_{min}^{-2}(\boldsymbol{z}).$
Then, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in direction.
Lastly, the resulting network f(;W^)f($\cdot$$\hat${W}) has the linear decision boundary defined by $\boldsymbol{\bar{w}}:=\frac{|J_{+}|}{\sqrt{m}}\boldsymbol{w}_{+}-\frac{|J_{-}|}{\sqrt{m}}\boldsymbol{w}_{-}$, i.e. $\mathrm{sign}(f(\boldsymbol{x}:\hat{W}))=\mathrm{sign}(\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle)$ for any $\boldsymbol{x}\in\mathbb{R}^{p}$.
We note that condition (i) goes beyond nearly orthogonal data regime studied in the existing literature since there is no upper bound to the magnitude of the signal $\boldsymbol{\mu}$, while condition (ii) contains the nearly orthogonal data regime studied in Frei et al. (2023b); Kou et al. (2023b). As discussed in Introduction, directional convergence of gradient descent in ReLU type neural networks was not established in the existing literature even under nearly orthogonal data regime. In addition, the characterizations of the convergent direction in Theorem 4.8 were obtained only for gradient flow trained on nearly orthogonal data by Frei et al. (2023b).
We show in Section 6 that these deterministic conditions on training data can be achieved with high probability under sufficient over-parametrization not only by (sG) but also by (PM).
The proof of Theorem 4.8 is given in Appendix B.